# Coincidences of Division Fields of an elliptic curve defined over a number field

### v2

@@ 5. Coincidence of division fields of two elliptic curves
Let $p$ be a prime and $r$ be the largest integer such that $\mathbb{Q}(\zeta_{p^{r}})\subseteq F\cap\mathbb{Q}(\zeta_{p^{\infty}})$. Suppose that $p>q$ for all primes $q\mid m$ and $v_{p}(n)>r$. Then, $v_{p}(n)=1$ (and $r=0$), unless $(m,p)=(2^{j},3)$ for some $j\geq 1$, in which case $r=0$ and $v_{p}(n)\leq 2$, or $v_{p}(n)=r+1$.
@@ Appendix A Derived groups of $\mathrm{GL}_{2}(m)$ and $\mathrm{SL}_{2}(m)$
By Diagram (5) and Proposition A.1, the image of $T$ in $\mathrm{SL}_{2}(m)$ generates the abelianization of $\mathrm{SL}_{2}(m)$, and its order divides $12$. Moreover, the image of $T$ in $\mathrm{SL}_{2}(m)$ has order $m$. Hence the abelianization of $\mathrm{SL}_{2}(m)$ has order dividing $\gcd(m,12)$. To prove that its order is exactly $\gcd(m,12)$, it suffices to prove it for $m=2,3$ and $4$, and then to use again Diagram (5). We define the commutators
@@
Let $G=\langle A,B\rangle\subseteq\mathrm{SL}_{2}(\mathbb{Z})$. For $m=2,3,4$, let $G_{m}$ be the image of $G$ is $\mathrm{SL}_{2}(3)$. By computing explicitely $G_{2}$, $G_{3}$ and $G_{4}$, we find that they have respectively order $3$, $8$ and $12$, so $\mathrm{SL}_{2}(m)/G_{m}$ has order $m$ and is abelian. Hence, as a subgroup of $\mathbb{Z}/12\mathbb{Z}$, it is isomorphic to $\mathbb{Z}/m\mathbb{Z}=\mathbb{Z}/\gcd(m,12)\mathbb{Z}$. ∎
@@
For an odd integer $m$, the derived group of $\mathrm{GL}_{2}(m)$ is $\mathrm{SL}_{2}(m)$, and so its abelianization is isomorphic to $(\mathbb{Z}/m\mathbb{Z})^{*}$. If $m$ is even, the derived group of $\mathrm{GL}_{2}(m)$ has index two in $\mathrm{SL}_{2}(m)$.

### v3

@@ 5. Coincidence of division fields of two elliptic curves
Let $p$ be a prime and $r$ be the largest integer such that $\mathbb{Q}(\zeta_{p^{r}})\subseteq F\cap\mathbb{Q}(\mu_{p^{\infty}})$. Suppose that $p>q$ for all primes $q\mid m$ and $v_{p}(n)>r$. Then, $v_{p}(n)=1$ (and $r=0$), unless $(m,p)=(2^{j},3)$ for some $j\geq 1$, in which case $r=0$ and $v_{p}(n)\leq 2$, or $v_{p}(n)=r+1$.
@@ Appendix A Derived groups of $\mathrm{GL}_{2}(m)$ and $\mathrm{SL}_{2}(m)$
By Diagram (5) and Proposition A.1, the image of $T$ in $\mathrm{SL}_{2}(m)$ generates the abelianization of $\mathrm{SL}_{2}(m)$, whose order divides $12$. Moreover, the image of $T$ in $\mathrm{SL}_{2}(m)$ has order $m$. Hence the abelianization of $\mathrm{SL}_{2}(m)$ has order dividing $\gcd(m,12)$. To prove that its order is exactly $\gcd(m,12)$, it suffices to prove it for $m=2,3$ and $4$, and then to use again Diagram (5). We define the commutators
@@
Let $G=\langle A,B\rangle\subseteq\mathrm{SL}_{2}(\mathbb{Z})$. For $m=2,3,4$, let $G_{m}$ be the image of $G$ is $\mathrm{SL}_{2}(m)$. By computing explicitely $G_{2}$, $G_{3}$ and $G_{4}$, we find that they have respectively order $3$, $8$ and $12$, and they are normal subgroups of $\mathrm{SL}_{2}(m)$. Then $\mathrm{SL}_{2}(m)/G_{m}$ is an abelian group of order $m$. Hence, as a subgroup of $\mathbb{Z}/12\mathbb{Z}$, it is isomorphic to $\mathbb{Z}/m\mathbb{Z}=\mathbb{Z}/\gcd(m,12)\mathbb{Z}$. ∎
@@
For an odd integer $m$, the derived group of $\mathrm{GL}_{2}(m)$ is $\mathrm{SL}_{2}(m)$, and so its abelianization is isomorphic to $(\mathbb{Z}/m\mathbb{Z})^{*}$. If $m$ is even, the derived group of $\mathrm{GL}_{2}(m)$ has index $2$ in $\mathrm{SL}_{2}(m)$.